Question 7 of 8 > At what temperature does 0.039925 moles of Ne in a 904.7 mL container exert a pressure of 0.964 atm?
Answer #1

Given:

P = 0.964 atm

V = 904.7 mL

= (904.7/1000) L

= 0.9047 L

n = 0.039925 mol

use:

P * V = n*R*T

0.964 atm * 0.9047 L = 0.0399 mol* 0.08206 atm.L/mol.K * T

T = 266 K

Answer: 266 K